Q. what is the period/fundamental period of 
|sinx| + |cosx|

Grade:11

1 Answers

Arun
6 years ago
If we consider  only sinx or cos x
Their fundamental period is 2\pi
Hence 
Period of sinx + cos x = 2\pi
But here actual question is
|Sin x|+|Cos x|
Mod is there hence graph will be on upper side of y axis
Hence their fundamental period is \pi